Mo can mow a lawn in 4 hours. His twin brother Joe can mow the same lawn in 5 hours. How long will it take Mo and Joe if they work together? Write your answer in hours and minutes
Answer by nerdybill(7384)

Mo can mow a lawn in 4 hours. His twin brother Joe can mow the same lawn in 5 hours. How long will it take Mo and Joe if they work together? Write your answer in hours and minutes
Let x = time (hours) it takes for both
then
x(1/4 + 1/5) = 1
multiplying both sides by 20:
x(5 + 4) = 20
x(9) = 20
x = 20/9
x = 2 hours and 13 minutes